OpenStudy (lopezking1):

Solve for x. 6 over x equals x over 12

OpenStudy (jack1):

6/x = x/12 \[\frac{ 6 }{ x } = \frac{ x }{ 12 }\] so cross multiply to get x^2 = ...something then solve for x

OpenStudy (lopezking1):

multiply both sides by 12.

OpenStudy (jack1):

yep, and then multiply both sides by x

OpenStudy (jack1):

that way u get rid of the 2 denominators (12 and x)

OpenStudy (lopezking1):

which simplifies to 72

OpenStudy (jack1):

on the LHS, yep so what does the RHS equal...?

OpenStudy (lopezking1):

x

OpenStudy (jack1):

nope, x^2 \[x ^{2}\]

OpenStudy (lopezking1):

so then square 72

OpenStudy (jack1):

times both sides by 12 6/x * 12 = x/12 * 12 72 / x = x then both sides by x 72 / x * x = x * x 72 = x^2 so x = sqrt 72 \[x = \sqrt{72}\]

OpenStudy (jack1):

yep, and sqrt 72 =...?

OpenStudy (lopezking1):

\[6\sqrt{2}\]

OpenStudy (jack1):

cool, i woulda gone with 8.5 ish, but yours is way more exact ;) understand it all tho dude?

OpenStudy (lopezking1):

Yeah thanks:D

OpenStudy (jack1):

no worries hey ;) best o luck